a)          Freddie has 6 toys cars and 3 toy buses, all different. 

i)              Freddie arranges these 9 toys in a line. Find the number of possible arrangements 

·           if there is a car at each end of the line and no buses are next to each other.



1
Expert's answer
2021-06-30T11:07:19-0400
"Car...Car...Car...Car...Car...Car"

Six cars we can arrange in "6!=720" ways.

Buses can be filled in the gap between two toy cars.

There are five gaps between the cars.

Three buses can be filled in five gaps in "P^5_3=\dfrac{5!}{(5-3)!}=60" ways.

So, the total number of ways to arrange is


"720(60)=43200"